### v1

# Automated Synthesis of Formally Verified Multi-Abstraction Function Summaries

Abstract
Function summaries, which characterize the behavior of code segments (typically functions) through preconditions and postconditions, are essential for understanding, reusing, and verifying software, particularly in safety-critical domains like aerospace embedded systems. However, these mission-critical legacy code serving as a valuable reused asset often lacks formal specifications. It is challenging to automatically generate function summaries for C programs, due to the existence of complex features such as loops, nested function calls, pointer aliasing, and so on. Moreover, function summaries should support multiple abstraction levels to meet diverse requirements, e.g. precise summaries capturing full functionality for formal verification and intuitive summaries for human understanding. To address these challenges, we first propose a novel framework that combines symbolic execution, large language models (LLMs), and formal verification to generate Relatively Strongest Postconditions (RSPs) and build function summaries that fully capture program behavior. Our approach leverages VST-A’s symbolic execution to precisely track program execution paths and state transitions, employs LLMs to infer loop invariants based on predefined templates, and uses Frama-C to guarantee soundness of generated summaries in an iterative refinement loop. Furthermore, from generated RSPs, we automatically synthesize strongest non-redundant postconditions expressed within given domain specific language. We compare our approach with existing work through extensive experiments.
Index Terms:
Function summary, Loop invariants, Symbolic execution, Large language models, Formal verification
I Introduction
In many safety-critical domains, such as aerospace, automotive, and medical systems, software often relies on legacy code that implements critical functionality but lacks formal specifications. This raises challenges for correct reuse of these codebases, particularly when such code is integrated into new safety-critical systems. A key requirement in these domains is the ability to automatically determine what a piece of code actually does, which we formalize as generating function summaries. These summaries are crucial for selecting correct implementations and ensuring their safe usage in safety-critical system design.
However, representing function summaries introduces additional challenges. For formal verification purpose, summaries must be precise enough to capture fine-grained functional behavior, with details necessary for rigorous proof construction. Yet, for human-centered tasks, e.g., selecting reusable components during design, overly detailed summaries make them difficult to understand for domain experts. Existing approaches fail to reconcile these needs: traditional verification tools require predefined specifications and thus produce property-specific postconditions to prove the given specifications, while design documentation suffers from ambiguity due to natural language description; and recent work on formal specification synthesis lacks multi-level abstractions to meet diverse requirements [1, 2, 3, 4].
This paper addresses these challenges by proposing a multi-granularity summary synthesis framework for C programs. We represent function summaries through pre-/postconditions, with an automatic precondition generation to derive the weakest sufficient initial conditions for correct execution (e.g. pointer validity). The key challenge lies in postcondition generation for C programs, which must account for complex features including loops, nested function calls, and complex pointer manipulations. They introduce fundamental difficulties: loops demand precise invariants to capture the behavior of unbounded iterations without predefined verification targets, complicating invariant synthesis; nested function calls introduce interprocedural control flow complexity; and pointer aliasing leads to state space complexity.
To address these challenges, we develop a combined approach which seamlessly integrate symbolic execution, formal verification and large language models (LLM) to generate relatively strongest postconditions (RSPs) of functions. Through symbolic execution, our tool can precisely track program execution paths and state changes, constructing strongest postconditions for straight-line code segments. For loops, we propose a template-guided LLM synthesis to generate invariants fitting constraint templates derived from symbolic execution data. Generated invariants are then iteratively refined via formal verification, with counterexamples fed back to the LLM until soundness is achieved. In the implementation, we leverage VST-A’s symbolic execution engine [5] for its ability to precisely track program states and path conditions at arbitrary execution points, combined with Frama-C [6] as our formal verifier to ensure summary correctness due to its automated verification and also ACSL-style assertions that are particularly amenable to LLM processing. Our framework tightly integrates VST-A, Frama-C, and LLMs, where each component’s strengths compensate for the others’ limitations.
Starting from the generated RSPs, we further develop an automated synthesis framework to generate more absract function summaries expressed in given domain specific languages (DSLs). Targeting aerospace software requirements as a case study, we design specialized DSLs focusing on different aspects such as side effects and numerical computation properties. Our framework is able to derive the strongest non-redundant postcondition expressible within each given DSL’s vocabulary, effectively distilling verification-grade RSPs into DSL summaries without losing essential information.
The rest of the paper is organized as follows: Sect. II introduces a motivating example and an overview of our approach. Sect. III and Sect. IV detail our core technical contributions on precise function summary generation and loop invariant generation. Sect. V further presents the DSL-based summary generalization. Sect. VI presents our implementation and Sect. VII evaluates our approach empirically. Finally, Sect. VIII presents the related work, and Sect. IX concludes with future directions. All materials related to tool implementation, LLM configurations, and experimental results are available at https://github.com/anon-hiktyq/ase2025-ARSPG.
II A Motivating Example And Our Approach
Our goal is to automatically generate function summaries at multiple abstraction levels, satisfying both rigorous verification and human-centric system design requirements. This task is particularly challenging for real-world code programs due to the existence of loops, nested function calls, and complex pointer manipulations. Traditional formal verification faces the same problem but it focuses exclusively on verifying whether a program satisfies a given specification rather than inferring complete summaries that fully characterize the program’s behavior. We illustrate these challenges using Fig. 1, which shows a code segment from the aerospace Sun Search software. TripleAbsMaxFun computes a sum-based threshold and the $L^{\infty}$ norm of an array in input ports, then conditionally writes this norm to a pointer in output port if it exceeds the threshold. It demonstrates the following challenges:
Figure 1: A Motivating Example
  • Multiple execution paths: The TripleAbsMaxFun function’s branching logic generates numerous distinct execution paths, each requiring analysis to ensure the summary captures all possible behaviors.
  • Pointer manipulations: The pointer parameter pIp introduces complexity due to memory aliasing and extensive side effects, e.g. the function modifies pIp->tmax and the memory location pointed to by pIp->ret, requiring the tracking of all these changes.
  • Loop and Function Call: The presence of a loop in getThreshold and the call to getThreshold within main add further complexity. Modeling loop behavior in terms of loop invariants becomes particularly challenging without predefined verification targets, as the absence of concrete proof objectives makes it difficult to assess and derive sufficiently strong invariants that fully capture the behavior of loop executions. Furthermore, nested function calls require precise interprocedural analysis to compose their effects in the calling context.
Although the precise summaries fully capture functional behavior including all state changes and side effects, their level of details can obscure high-level design intention. For instance, the TripleAbsMaxFun function’s core logic of computing the maximum absolute value of three inputs becomes buried under low-level operational summaries. This reveals the need for verified yet human-interpretable summaries expressed in domain specific formalisms.
II-A Our Approach
Our approach to generating program summaries is structured into a two-phase process: summary generation and summary generalization, as shown in Fig. 2.
The summary generation phase employs mainly symbolic execution to automatically derive precise, first-order logic summaries from input functions by exploring all execution paths. Two program constructs require specialized handling:
  • Loop invariant generation: When a loop is encountered during symbolic execution, a loop invariant is required to summarize its behavior without unfolding the loop many times for the symbolic execution to proceed. We develop a key module that combines symbolic execution, LLMs and formal verification to generate precise and sound loop invariants. Especially, using symbolic execution results and constraint templates, it produces precise invariants able to characterize loop behavior without relying on predefined verification targets.
  • Callee summary generation: During symbolic execution, function calls are handled by generating reusable callee summaries. These are instantiated with current symbolic states to derive the function call postconditions. For nested calls, a call stack is introduced to track the execution contexts, constructing summaries bottom-up for composition.
Ideally, once symbolic execution is completed for a function, the disjunction of assertions specifying all reachable paths and final symbolic states constitutes the relative strongest postcondition (RSP) of the function, which serves as its most precise summary.
The specification generalization phase takes the precise function summary derived from the first phase and transforms it into a more generalized and abstract representation within a given domain-specific language (DSL). As illustrated in Fig. 2, we implement an algorithm to find the best conjunction in DSL from the initial function summary, which is analogous to computing a strongest non-redundant clause set. It involves simplifying the complex disjunctive precise summary into a conjunction of clauses expressible in the chosen DSL.
Figure 2: Overview of Our Architecture
III Precise Summary Generation
III-A Symbolic Representation
We use a flattened memory model to represent symbolic states, explicitly tracking pointer aliasing and memory dependencies. Before giving the definition, we introduce some notations. A program point $p$ is a distinct location in a program’s control flow, with an associated scope determining accessible variables. Within a given scope, the memory is modeled as a collection of atomic memory locations $\mathcal{L}_{p}$, representing flattened program variables and addressable memory units. Each $v\in\mathcal{L}_{p}$ is a unique, indivisible unit storing a value of a base type, e.g., int, pointer address etc.
A symbolic state $\sigma_{p}$ at program point $p$ maps $\mathcal{L}_{p}$ to their respect symbolic values in $\mathcal{V}$. Our memory model assumes that at any program point $p$, every atomic memory location in $\mathcal{L}_{p}$ is accessible and initialized. The symbolic execution of a program is then a sequence of state transitions, staring from the initial state $\sigma_{p_{0}}$ at initial program point $p_{0}$.
Limitations: Building upon the flattened memory model, we restrict our analysis to C programs with non-recursive data structures and non-recursive function calls. This ensures all memory locations can be statically determined and finitely represented in $\mathcal{L}_{p}$.
III-B Symbolic Execution
Symbolic execution is a program analysis technique that explores program execution paths using symbolic values instead of concrete inputs. Given a function $\mathcal{F}$, its symbolic execution starts with a precondition Pre, defined as $PC_{0}\land\mathrm{StateForm}(\sigma_{p0})$, where $\sigma_{p0}$ is the initial state mapping inputs to initial symbolic values, $\mathrm{StateForm}(\sigma_{p0})$ is the logical formula, i.e. the conjunction of equalities describing the values of all memory locations in $\sigma_{p0}$ in terms of initial symbolic values, and $PC_{0}$ the constraints on these symbolic values. By default, Pre ensures that all memory locations of $\sigma_{p0}$ are accessible and initialized. An execution path $\pi$ tracks symbolic values of all atomic memory locations, accumulating branch constraints. Upon completion, it yields $(PC_{\pi},\sigma_{p,\pi})$, where $\sigma_{p,\pi}$ is the final symbolic state and $PC_{\pi}$ is the path condition for path $\pi$ to be executable.
We define two procedures: $\mathrm{SEStart}(\mathcal{F},\textit{Pre},p)$ performs symbolic execution from any program point $p$ for function $\mathcal{F}$ with precondition Pre, till a loop or a function call is reached, and $\mathrm{SECollect}(\mathcal{F},\textit{Pre},p)$ collects the symbolic results composed of sets of (path, state) for program point $p$, i.e.
$\left\{(PC_{\pi},\sigma_{p,\pi})\left|\begin{array}[]{l}\pi\text{is an execution path}\\ \text{reachable from }p_{0}\text{ to }p\end{array}\right.\right\}$
Denote the function’s exit point by $p_{-1}$, then $\mathrm{SymExecCollect}(\mathcal{F},\textit{Pre},p_{-1})$ provides all accumulated path conditions and states at termination of $\mathcal{F}$.
Loops and function calls pose challenges for complete path exploration. To address these constructs, our approach integrates symbolic execution with LLMs and formal verification to generate loop invariants to over-approximate iterative behavior and callee function summaries to handle (nested) function calls. We detail these techniques in Sections III-D and IV respectively.”
III-C Relative Strongest Postconditions
Performing symbolic execution on a program segment from its entry $p_{0}$ to any program point $p$ yields a set of pairs $(PC_{\pi},\sigma_{p,\pi})$ for all reachable paths $\pi$ targeting at $p$. The relative strongest postcondition (RSP) of $\mathcal{F}$ at $p$ with respect to Pre, denoted by $\textit{RSP}(\mathcal{F},\textit{Pre},p)$, is a logical formula describing the disjunction of these (path, state) pairs of $p$ reachable from $p_{0}$ satisfying Pre, defined as blow:
$PC_{0}\land\Bigl{(}\bigvee_{i=1}^{m}\bigl{(}PC_{\pi_{i}}\land\mathrm{StateForm}(\sigma_{p,\pi_{i}})\bigr{)}\Bigr{)}$ (1)
where $\mathrm{StateForm}(\sigma_{p,\pi_{i}})$ is the logical formula specifying the symbolic state $\sigma_{p,\pi_{i}}$ in terms of initial symbolic values recorded by $PC_{0}$. Specially, $\textit{RSP}(\mathcal{F},\textit{Pre},p_{-1})$ for the exit point $p_{-1}$ is the final RSP of $\mathcal{F}$ with respect to Pre at termination.
The ”relative” in RSP reflects its dependence on loop invariant accuracy during symbolic execution. Perfect invariants yield the true strongest postcondition, but in practice, we usually obtain over-approximation of the actual reachable state space of loops.
III-D Function Summary Generation
Our approach automatically derives RSPs from function preconditions through path-sensitive symbolic execution, producing formally verifiable function summaries. Algorithm 1 provides the complete realization of this process.
Lines 1-5 initialize the default precondition (setting up initial values for all atomic memory locations), loop and function call stacks, and function annotations. Symbolic execution of function $\mathcal{F}$ then begins from its entry point $p_{0}$ with the specified precondition by invoking $\mathrm{SEStart}$ (Line 6), which will terminate whenever a loop or a function call is reached.
Upon encountering a loop or a function call at program point $p$, we construct the precondition from the symbolic execution results obtained at $p$ via $\mathrm{SECollect}$ (Lines 7-9). If it is a loop (Lines 10-11), $\mathrm{LoopInvGen}$ is invoked to synthesize a loop invariant and returns an updated function $\mathcal{F}$ that annotates the loop with its invariant. If it is a function call (Lines 12-13), a recursive $\mathrm{CalleeSum}$ procedure is invoked to obtain the callee function summary and returns an updated function $\mathcal{F}$ that annotates the function call with the callee summary. For both of them, the symbolic execution can re-proceed with the corresponding annotations (Line 14).
After exploring all paths to the function’s exit, the algorithm collects accumulated path conditions and their corresponding symbolic states. These are then combined with the function’s initial precondition to formally construct the RSP of $\mathcal{F}$ (Lines 15-16). Finally, this inferred summary is integrated with the function and returned (Lines 17-18).
Input: Function $\mathcal{F}$
Output: Function $\mathcal{F}$ with summary $\mathcal{S}$
1
2$(\mathcal{F},PC_{0},\sigma_{p_{0}})\leftarrow\mathrm{DefaultAnnotation}(\mathcal{F})$;
3 $\textit{Pre}\leftarrow(PC_{0}\land\mathrm{StateForm}(\sigma_{p_{0}}))$;
4 $\textit{lS}\leftarrow\mathrm{InitLoopStack}()$;
5 $\textit{cS}\leftarrow\mathrm{InitCallStack}()$;
6 $\textit{aF}\leftarrow\mathrm{InitAnnoFuncs}()$;
7
[…]
9
10while $p\neq p_{-1}$ and SE terminated do
11
12      $\left\{\left(PC_{\pi_{i}},\,\sigma_{p,\pi_{i}}\right)\right\}_{i=1}^{m}\leftarrow\mathrm{SECollect}(\mathcal{F},\textit{Pre},p)$;
13       $P\leftarrow\textit{RSP}(\mathcal{F},\textit{Pre},p)$;
14
      // only outer loop entry can be reached
15       if $p=p_{\textit{loop}}$ then
             // P is the precondition of the loop
16             $\mathcal{F}\leftarrow$ $\mathrm{LoopInvGen}(\mathcal{F},\,p,\,\textit{lS},\,\textit{cS},$ $\,\textit{aF},\,P,\,\textit{LLM}$);
17
18      if $p=p_{\textit{callee}}$ then
19             $\mathcal{F}_{\textit{callee}}\leftarrow\mathrm{CalleeSum}(\mathcal{F}_{\textit{callee}},\,\textit{lS},\,\textit{aF})$;
20
21      $\mathrm{SEStart}(\mathcal{F},\textit{Pre},p)$;
22
23
24$\left\{\left(PC_{\pi_{i}},\,\sigma_{p_{-1},\pi_{i}}\right)\right\}_{i=1}^{m}\leftarrow\mathrm{SECollect}(\mathcal{F},\textit{Pre},p_{-1})$ ;
25
[…]
27
28$\mathcal{F}\leftarrow(\mathcal{F},\mathcal{S})$;
29
30return $\mathcal{F}$;
Algorithm 1 Function Summary Generation $\mathrm{FunSum}$
III-E Callee Summary Generation
Algorithm 2 enables a compositional approach to handle nested function calls by first deriving, then leveraging summaries of individual functions within a call graph. As shown in Algorithm 2, if the callee function already has a generated summary (i.e., it’s in aF), its summary is immediately returned. Otherwise, the $\mathrm{FunSum^{\prime}}$ algorithm is invoked to generate the summary for the callee function (Note: $\mathrm{FunSum^{\prime}}$ differs from $\mathrm{FunSum}$ only by omitting the call stack initialization in Line 4). This newly derived summary is then stored and annotated to the function, making it available for any future calling context.
[…]
Output: Function $\mathcal{F}$ with summary $\mathcal{S}$
1
2if $\mathcal{F}\in\textit{aF}$ then
3       return
[…]
5 $\mathrm{FunSum^{\prime}}(\mathcal{F})$;
[…]
7return $\mathcal{F}$;
Algorithm 2 Callee Summary Generation $\mathrm{CalleeSum}$
IV Loop Invariant Generation
The general loop invariant generation problem can be defined as follows:
Definition 1 (Loop invariant generation problem).
Given a loop $l:\!$ while B do C of program $\mathcal{P}$, a precondition $P$ that holds at entry $l$, and an (optional) postcondition $Q$, the loop invariant generation problem is to find a logical formula Inv that satisfies the following conditions:
  • Base: $P\Rightarrow\textit{Inv}$, i.e., the invariant Inv holds initially;
  • Preservation: $\{\textit{Inv}\land B\}\ C\ \{\textit{Inv}\}$, i.e. if Inv holds at the start of a loop iteration and loop condition $B$ is true, executing loop body $C$ guarantees Inv remains true;
  • Termination: $\textit{Inv}\land\neg B\Rightarrow Q$, i.e. upon loop termination, Inv remains true and implies postcondition $Q$.
As defined above, the conventional invariant generation problem requires a pregiven postcondition to synthesize sufficiently strong invariants that imply the postcondition. Our invariant generation algorithm, to be given below, operates without postcondition guidance, yet produces invariants of comparable strength. By combining symbolic execution with LLM-based template instantiation, it derives comprehensive and precise invariants for non-nested loops. Sect. VII-A experimentally validates this capability against standard benchmarks.
[…]
Output: Function $\mathcal{F}$ with loop invariant $\mathcal{I}$ for loop $\mathcal{F}_{\textit{loop}}$
1
// $\mathcal{F}_{\textit{loop}}$ is loop body at $p_{\textit{loop}}$ in $\mathcal{F}$
2 $\textit{lS}.push(\mathcal{F}_{\textit{loop}})$;
[…]
4
[…]
// is the loop Outer Loop? always yes
6 $\mathrm{Assert}(\textit{lS}.empty())$;
7
[…]
9
[…]
11
12$\mathcal{F}\leftarrow((\mathcal{F}-\mathcal{F}_{\textit{loop}}),\,(\mathcal{F}_{\textit{loop}},\mathcal{I}))$ ;
13
14while $\neg\mathrm{Verify}(\mathcal{F})$ do
15       $\mathcal{F}\leftarrow\mathrm{LoopInvRefinement}(\mathcal{F},\textit{LLM})$;
16
17
18return $\mathcal{F}$;
Algorithm 3 Loop Invariant Generation $\mathrm{LoopInvGen}$
Algorithm 3 presents a three-phase invariant generation algorithm for an outermost loop ($F_{loop}$) and any loops nested within its body. Initially (Lines 1-4), the target loop is pushed to the loop stack for handling, then $\mathrm{LoopAnalysis}$ is invoked to generate invariants for inner nested loops and annotate them in $F_{loop}$, after which the target loop is popped; Subsequently (Lines 5-7), the invariant templates are generated via $\mathrm{LoopInvTemGen}$, then the template-guided invariant synthesis via LLM is performed by invoking $\mathrm{OuterLoopLLMGen}$, followed by the annotation of the invariants of $F_{loop}$ to the main function; Lastly (Lines 8-9), $\mathrm{LoopInvRefinement}$ performs fine-grained verification-driven refinement, where the LLM interacts with the formal verifier iteratively to adjust unverified invariants until a formally verified one is achieved.
[…]
1
[…]
3 $(\mathcal{F^{\prime}},PC_{0}^{\prime},\sigma_{p_{0}}^{\prime})\leftarrow\mathrm{DefaultAnnotation}(\mathcal{F^{\prime}})$;
4 $\textit{Pre}\leftarrow(PC_{0}^{\prime}\land\mathrm{StateForm}(\sigma_{p_{0}^{\prime}}))$;
5
[…]
7
8while $p\neq p_{-1^{\prime}}$ do
9       if $p=p_{\textit{loopEntry}}$ then
10             $\textit{lS}.push(\mathcal{F}_{\textit{loop}})$;
11
12      if $p=p_{\textit{loopExit}}$ then
13             $\mathcal{F}_{\textit{loop}}\leftarrow\textit{lS}.pop()$;
             // call LLM
14             $\mathcal{F}\leftarrow$ $\mathrm{NestedLoopLLMCall}(\mathcal{F},\mathcal{F}_{\textit{loop}},\textit{LLM})$;
15
16      if $p=p_{\textit{callee}}$ and SE terminated then
17             $\mathcal{F}_{\textit{callee}}\leftarrow$ $\mathrm{CalleeSum}(\mathcal{F}_{\textit{callee}},\,\textit{cS},\,\textit{aF})$;
18             $\mathrm{SEStart}(\mathcal{F^{\prime}},\textit{Pre},p)$;
19
20
21
22$\left\{\left(PC_{\pi_{i}},\,\sigma_{p_{-1^{\prime}},\pi_{i}}\right)\right\}_{i=1}^{m}\leftarrow\mathrm{SECollect}(\mathcal{F^{\prime}},\textit{Pre},p_{-1^{\prime}})$ ;
23
[…]
25
[…]
27
[…]
29
30return $\mathcal{F},\textit{uV},\textit{nV}$ ;
Algorithm 4 loop Analysis $\mathrm{LoopAnalysis}$
IV-A Loop Analysis
Algorithm 4 implements the first phase, combining symbolic execution and LLMs to generate invariants for nested loops in $\mathcal{F}$ and identify unchanged/non-inductive variables. It begins by unfolding $\mathcal{F}$ and initiating symbolic execution at the loop body start (Lines 1-4). During execution, it handles three cases: pushing encountered nested loops to the stack at loop entries (Lines 6-7), popping completed loops and annotating them with invariants through the LLM-based invariant synthesis $\mathrm{NestedLoopLLMCall}$ (Lines 8-10), and handling symbolic execution of function calls via $\mathrm{CalleeSum}$ (Lines 11-13). The phase concludes by producing $\mathcal{F}$’s loop body RSP, and identifying the sets of unchanged and non-inductive variables (Lines 14-18).
IV-B Template Generation
The above $\mathrm{LoopAnalysis}$ phase extracts symbolic execution data from loop unfoldings to provide inputs for $\mathrm{LoopInvTemGen}$, which then generates a set of invariant templates. These templates have parameters partially determined by symbolic execution results, while the remaining undetermined parameters are completed using the subsequent LLM synthesis method (as presented in Algorithm 3). Below we introduce four forms of invariant templates.
  • Unchanged variables: For any variable $v\in uV$ and any execution path $\pi_{j}$ to the loop entry $p$, the invariant component $\mathcal{I}_{v,\pi_{j}}$ is:
    $\mathcal{I}_{v,\pi_{j}}\equiv PC_{\pi_{j}}\Rightarrow v=\sigma_{p,\pi_{j}}(v)$
    asserting that, if execution follows path $\pi_{j}$, $v$ consistently retains its symbolic value $\sigma_{p,\pi_{j}}(v)$ at the loop entry.
  • Loop skip conditional property: For a path $\pi_{j}$ to $p$, and a loop condition $B$ where $v\in\mathcal{L}_{p}$, if the condition $B$ for path $\pi_{j}$ is false, the loop is skipped without altering variables. The invariant component $\mathcal{I}_{v,\pi_{j}}$ is:
    $B^{\prime}\leftarrow B\land(PC_{\pi_{j}}\land\mathrm{StateForm}(\sigma_{p,\pi_{j}}))$
    $\mathcal{I}_{v,\pi_{j}}\equiv\neg B^{\prime}\Rightarrow v=\sigma_{p,\pi_{j}}(v)$
    This ensures $v$ retains its initial value. Both the above two types of invariants satisfy the base and preservation conditions.
  • Non-inductive variables: For a non-inductive variable $v$ (typically overwritten in the first iteration) and any path $\pi_{j}$ to $p$:
    $B^{\prime}\leftarrow B\land(PC_{\pi_{j}}\land\mathrm{StateForm}(\sigma_{p,\pi_{j}}))$
    $\mathcal{I}_{v,\pi_{j}}\equiv(B^{\prime}\Rightarrow((v=\sigma_{p,\pi_{j}}(v))\lor\Phi(v,\mathcal{L}_{p}\cup\mathcal{V})))$
    where $\Phi(v,\mathcal{L}_{p}\cup\mathcal{V})$ is the synthesized inductive invariant.
  • Inductive variables: For an inductive variable $v$ (which is neither unchanged nor non-inductive and whose value relates to its previous value after the iteration) and any path $\pi_{j}$ to $p$:
    $B^{\prime}\leftarrow B\land(PC_{\pi_{j}}\land\mathrm{StateForm}(\sigma_{p,\pi_{j}}))$
    $\mathcal{I}_{v,\pi_{j}}\equiv(B^{\prime}\Rightarrow\Phi(v,\mathcal{L}_{p}\cup\mathcal{V}))$
This template-driven approach is a cornerstone of our loop invariant geneartion algorithm. By using precise symbolic execution, it simplifies the LLM’s task, shifting from full invariant generation to completing template invariant components. It also provides inherent syntactic guidance, easing the LLM’s adherence to formal logical syntax. This approach guarantees strong invariants that capture all relevant variable behaviors across every execution path, thereby closely over-approximating the set of all possible variable assignments within the loop.
IV-C Loop Invariant Refinement
Loop invariants generated by the LLM may not always be sound or sufficiently strong. To ensure their correctness, we implement a fine-grained, iterative refinement framework. This system leverages verifier errors and the LLM’s self-correction to continuously refine invariants. Verification failures fall into four categories, each prompting a specific action:
  • Syntax error: Verifier reports a parse error and the LLM repairs the syntax error directly.
  • Invariant failure (postcondition passed): Invariants fails Base or Preservation condition, yet surprisingly implies the post-condition (if given). The LLM adjusts the invariant to correct the failed property.
  • Invariants correct (postcondition failed): Invariant satisfies Base and Preservation, but is too weak, failing Termination postcondition (if given). The LLM strengthens the invariant.
  • Fundamental failure: Invariant fails Base, Preservation, and Termination. A complete re-generation by the LLM is triggered.
If errors persist after multiple repair rounds, we apply a final filtering step to iteratively prune any invariant components consistently failing.
V Summary Generalization
The RSPs generated in the previous stage are crucial for verification of safety-critical systems, however, they can be overly detailed for human understanding. To address this, starting from the RSPs, we propose a summary generalization procedure to generate summaries expressed within given Domain-Specific Languages (DSL). This process distills implementation details while guarantees to derive the strongest postconditions with respect to the DSL.
In our approach, the DSL is constructed by manually providing parameterized clauses, called candidate clauses, which serve as reusable templates for abstract summary generation. These clauses are then instantiated using parameters drawn from the RSP’s symbolic variables and constants, thereby aligning the DSL clause set with the RSP’s context.
Our objective is to identify the optimal DSL conjunction expressed within the chosen DSL, given the generated RSP. This is achieved by first instantiating the DSL template formulas to obtain a clause set and then finding the conjunction of the strongest non-redundant clause set with respect to the given RSP. Let $C$ be the instantiated clause set, then:
Definition 2 (Strongest non-redundant clause set).
A set of clauses $D\subseteq C$ is a strongest non-redundant clause Set of $C$ with respect to the RSP $Q$, if it satisfies:
  1. 1.
    Entailment: For every $c\in D$, $Q\models c$.
  2. 2.
    Non-Redundancy: For every $c\in D$, $D\setminus\{c\}\not\models c$.
  3. 3.
    Maximality: There is no $D^{\prime}\subseteq C$ such that $D\subset D^{\prime}$ and $D^{\prime}$ satisfies (1) and (2).
Algorithm 5 computes the strongest non-redundant set for a given RSP $Q$ and a set of parameterized DSL candidate formulas $\Omega$. It operates in three phases: First (Line 1), function $\mathrm{Instantiate}$ generates instantiated clauses $C$ from parameterized $\Omega$ using symbolic variables $\mathcal{L}_{-1}$ and constants $\mathcal{V}$ from RSP $Q$; Next (Lines 3-5), it filters $C$ that are implied by the RSP $Q$, resulting in the initial set $E$; Third (Lines 7-16), a refinement process is performed: each clause $c_{\textit{new}}$ from $E$ is added to the resulted set $D$ only if it is not already entailed by $D$. This step also ensures non-redundancy by removing any existing clause $d_{\textit{old}}$ from $D$ that is entailed by $c_{\textit{new}}$.
Input : RSP $Q$ of function $\mathcal{F}$, parameterized DSL $\Omega$, symbolic variables $\mathcal{L}_{-1}$ at $\mathcal{F}$’s return and their initial values $\mathcal{V}$
Output : $D$
1 $C\leftarrow\mathrm{Instantiate}(\Omega,\mathcal{L}_{-1},\mathcal{V})$;
2 $E\leftarrow\emptyset$;
3 foreach $c\in C$ do
4       if $\mathrm{Entails}(Q,c)$) then
5             $E\leftarrow E\cup\{c\}$;
6
7
[…]
9 foreach $c_{\textit{new}}\in E$ do
10       if $\mathrm{Entails}(D,c_{\textit{new}})$ then
11             continue;
12
13      $D_{\textit{to\_remove}}\leftarrow\emptyset$;
14       foreach $d_{\textit{old}}\in D$ do
15             if $\mathrm{Entails}(\{c_{\textit{new}}\},d_{\textit{old}}$) then
16                   $D_{\textit{to\_remove}}\leftarrow D_{\textit{to\_remove}}\cup\{d_{\textit{old}}\}$;
17
18
19      foreach $d_{\textit{rem}}\in D_{\textit{to\_remove}}$ do
20             $D\leftarrow D\setminus\{d_{\textit{rem}}\}$;
21
22      $D\leftarrow D\cup\{c_{\textit{new}}\}$;
23
24return $D$;
Algorithm 5 Compute Strongest Non‑Redundant Set
We have proved the correctness of this algorithm, as stated by the following theorem:
Theorem 1.
Algorithm 5 returns a strongest non-redundant clause set (i.e. $D$) of $C$ with respect to $Q$, i.e. $D$ satisfies the Entailment, Non-redundancy and Maximality conditions of Def. 2.
The termination of the algorithm is ensured by the finite nature of the DSL templates and the memory locations. Computation in our implementation is accelerated through several optimizations. We cache intermediate entailment results to eliminate redundant checks in the Entails phase. Additionally, heuristic sorting prioritizes potentially stronger clauses in $C$, and inductive entailment with concrete integers, informed by clause templates, effectively prunes the search space.
VI Implementation
VI-A Toolchain Integration of VST-A, Frama-C, and LLM
In our implementation, we combine the strengths of symbolic execution, large language models (LLMs), and formal verification to achieve precise, efficient and sound function summary generation. We use VST-A [5] as our foundational symbolic execution engine, which generates precise path conditions and symbolic states for straightline code through Coq-verified strongest postcondition derivation. However, VST-A requires manual annotation for loop invariants and callee function summaries, and its output may include unproven lemmas requiring manual Coq proofs. To address these limitations, we leverage LLMs to automatically synthesize candidate invariants, and meanwhile employ Frama-C [6] as our formal verifier to ensure soundness and automate verification. The integration of Frama-C with SMT solvers enables automatic verification, while its ACSL (ANSI/ISO C Specification Language) assertion language offers LLM-friendly syntax to facilitate iterative collaboration. Through collaboration, Frama-C refines incorrect invariants learned from LLM through counterexamples and guides LLMs to finally obtain sound invariants.
VI-B Implementation Details
We detail the realization of some key functions here. Function $\mathrm{DefaultAnnotation}$ establishes the initial precondition of a given function for its symbolic execution to start. While manual setup is possible, our tool automatically provides VST-A a default precondition for each function, which specifies the validity assumptions over atomic memory locations within the initial program point’s scope.
For LLM-based invariant synthesis, Both functions $\mathrm{OuterLoopLLMCall}$ and $\mathrm{NestedLoopLLMCall}$ synthesize loop invariants by integrating LLMs with a zero-shot, Chain-of-Thought (CoT) prompting approach. CoT improves LLM understanding via contextual questions before invariant generation.
A key feature of our toolchain is the interoperability between VST-A and Frama-C, despite their different assertion languages. During the $\mathrm{LoopInvGen}$ phase, LLM-generated loop invariants in ACSL are translated into VST user assertion format. This enables VST-A to correctly interpret and leverage them for symbolic execution across loop and function boundaries. After symbolic execution, VST-A’s generated RSPs are converted from its assertion format to ACSL for subsequent Frama-C automated verification. To achieve these, we establish an automatic translation that enables the conversion of assertions in both directions (see the supplementary material).
VII Experiments
In this section, we refer to our approach for automatic relative strongest postcondition generation as ARSPG.
VII-A Experiments on Loop Invariant Generation
VII-A1 Performance on benchmarks
We evaluated our loop invariant generation tool on three distinct benchmarks:
  • syGus [7]: Simpler, single-loop structures, for fundamental invariant generation.
  • OOPSLA [8]: More challenging, multi-loop and nested loops, for robustness.
  • SV-COMP: Adopted from AutoSpec’s SV-COMP22 selection [9], also featuring multi-loop and nested loops.
Our LLM selection balanced model maturity, prevalence, and cutting-edge performance:
  • GPT-4O-mini and GPT-3.5-turbo: Smaller, established models for efficiency and baseline.
  • GPT-4O and Deepseek-v3: Current, widely adopted models for general and advanced code comprehension.
  • Claude-3.7-sonnet: Strongest model for code understanding, setting peak performance.
We measured correctness via three key metrics:
  • Syntax rate (Syn.): Percentage of syntactically correct invariants.
  • Valid rate (Val.): Percentage of syntactically correct invariants satisfying Base and Preservation.
  • Accuracy rate (Acc.): Percentage of syntactically correct invariants satisfying Base, Preservation, and Termination.
These metrics are evaluated at different Pass@k, indicating that the system may attempt multiple generations until a valid invariant is found.
TABLE I: Loop Invariant Generation Performance Overview
Benchmark Model Pass@1 Pass@3 Pass@5
Syn. Val. Acc. Syn. Val. Acc. Syn. Val. Acc.
syGas GPT-3.5-turbo 94.74% 82.71% 69.92% 100.00% 93.98% 82.71% 100.00% 96.24% 87.22%
syGas GPT-4O-mini 89.37% 78.95% 65.41% 100.00% 91.73% 81.20% 100.00% 95.49% 85.71%
syGas GPT-4O 98.50% 93.98% 87.22% 100.00% 96.99% 94.74% 100.00% 99.25% 96.99%
syGas DeepSeek-V3 98.50% 94.74% 87.97% 100.00% 99.25% 98.50% 100.00% 99.25% 98.50%
syGas Claude-3.7-sonnet 100.00% 99.25% 97.74% 100.00% 100.00% 99.25% 100.00% 100.00% 99.25%
OOSPLA GPT-3.5-turbo 82.61% 54.35% 41.30% 86.96% 65.22% 45.65% 86.96% 67.39% 47.83%
OOSPLA GPT-4O-mini 89.37% 78.95% 65.41% 89.13% 86.96% 43.48% 89.13% 86.96% 47.83%
OOSPLA GPT-4O 91.30% 69.57% 60.87% 95.65% 76.09% 73.91% 100.00% 82.61% 76.08%
OOSPLA DeepSeek-V3 91.30% 67.39% 63.04% 97.83% 82.61% 78.26% 97.83% 86.96% 82.60%
OOSPLA Claude-3.7-sonnet 93.48% 69.67% 63.04% 97.83% 80.43% 76.09% 97.83% 93.57% 86.96%
SV-COMP GPT-3.5-turbo 52.38% 52.38% 23.81% 61.90% 61.90% 38.10% 76.19% 76.19% 52.38%
SV-COMP GPT-4O-mini 61.90% 61.90% 28.57% 80.95% 80.95% 42.86% 85.71% 80.95% 57.14%
SV-COMP GPT-4O 90.48% 90.48% 52.38% 100.00% 100.00% 71.43% 100.00% 100.00% 76.19%
SV-COMP DeepSeek-V3 80.95% 76.19% 47.62% 95.24% 90.48% 66.67% 95.24% 90.48% 71.43%
SV-COMP Claude-3.7-sonnet 100.00% 100.00% 71.43% 100.00% 100.00% 71.43% 100.00% 100.00% 76.19%
Table I lists the performance results of our approach ARSPG on the three benchmarks. ARSPG notably excels at handling single-loop numerical programs (as seen in sysGus), and can also generate accurate invariants for multi-loop programs, despite their complexity. Moreover, these results show that both the use of stronger LLM models and consistent improvement across Pass@k significantly benefit performance on these more challenging problems.
Discussion
We notice that the accuracy on OOPSLA and SV-COMP benchmarks declined compared to sysGus. This is due to two reasons:
  • For nested loops, symbolic execution struggles to provide necessary loop preconditions for LLMs to generate effective invariants.
  • Many programs, especially in OOPSLA, lack explicit or easily extractable loop conditions, impeding precise invariant inference.
Despite these hurdles, our tool achieved notable accuracy on both OOPSLA and SV-COMP. Consistent improvement from Pass@1 to Pass@5 underscores its potential and robustness in complex multi-loop scenarios.
VII-A2 Comparison with existing work
TABLE II: Loop Invariant Generation Performance Comparison
Method syGas OOSPLA SV-COMP
CODE2INV 54.9%
HOLA 93.5%
CLN2INV 93.2%
AutoSpec 85.7% 82.6% 76.2%
ACLInv 79.9% 76.1% 76.2%
ARSPG 97.0% 76.1% 76.2%
ARSPG-Best 99.3% 87.0% 76.2%
In Table II, we evaluated ARSPG under Pass@1 and zero-shot conditions against a range of existing tools for loop invariant generation, broadly categorized into traditional and LLM-based approaches. Traditional tools like CODE2INV[10], CLN2INV[11], and HOLA[8] are typically optimized for specific, single datasets. LLM-based tools such as AutoSpec and ACInv provided a benchmark, for which we compared to their best reported results.
ARSPG was assessed in two configurations: ARSPG, using the GPT-4O model with a Pass@5 setting, and ARSPG-Best, representing optimal performance with the Claude-3.7-sonnet model at Pass@5. As presented in Table II, when ARSPG is augmented with advanced LLMs, it outperforms CODE2INV and CLN2INV on the syGus benchmark, with ARSPG-Best achieving a remarkable 99.3%. On the more challenging OOPSLA benchmark, ARSPG-Best also demonstrates strong performance at 87.0%, outperforming both AutoSpec and ACInv. For the SV-COMP benchmark, our results are consistent with those of AutoSpec and ACInv, achieving 76.2%. These findings collectively underscore the effectiveness of our approach in generating high-quality loop invariants across diverse benchmarks.
To demonstrate ARSPG’s capability on invariant generation in absence of verification targets, we conducted experiments on syGus benchmarks by deliberately masking loop postconditions. Table IV shows the results: With masked postconditions, ARSPG achieves 80.54% verification accuracy (vs. 87.22% unmasked), retaining high invariant precision despite missing targets. This confirms our method’s robustness when termination conditions are unavailable.
VII-A3 Effectiveness
TABLE III: Loop Invariant Generation Effectiveness: Ablation Study
Method Pass@1 Pass@3
Syn. Val. Acc. Syn. Val. Acc.
Only LLM 86.25% 23.75% 18.75% 97.50% 37.50% 28.75%
LLM with SE 90.00% 47.50% 32.50% 100.00% 67.50% 52.50%
ARSPG 100.00% 100.00% 100.00% 100.00% 100.00% 100.00%
To evaluate the effectiveness of different techniques applied in our tool, we conducted an ablation study on 80 selected loop programs. We compared our full approach against two distinct baselines: Only LLM employs a LLM for zero-shot invariant generation without external guidance; LLM with SE is augmented with symbolic execution to guide the LLM’s generation via templates and symbolic execution information.
As demonstrated in Table III, ARSPG consistently outperformed both only LLM and LLM with SE, especially for valid rate and accuracy rate. Notably, integrating symbolic execution with LLMs already yields great improvements over pure LLM baselines, through template-based synthesis under same LLM call time.
TABLE IV: Invariant Generation With/Without Postcondition Guidance
GPT-4O Pass@1 Syn. Val. Acc.
Masked ARSPG 96.99% 96.24% 80.54%
ARSPG 98.50% 93.98% 87.22%
To sum up, the advantage of our approach includes: First, through a novel synthesis mechanism integrating symbolic execution with template-based LLM, ARSPG can still generate precise invariants without relying on termination conditions (see Table IV); Second, ARSPG prevents overfitting to fixed datasets through generic invariant template and flexible LLM call, demonstrating consistent performance across diverse programs; Third, ARSPG’s zero-shot approach for LLM-based invariant generation eliminates example induced bias and is able to adapt to new invariant generation tasks without retraining or prompt engineering.
VII-B Function Summary Generation Performance
We evaluated ARSPG for function summary generation on 51 programs from AutoSpec [1] benchmark and successfully verified 29 programs under Pass@1 zero-shot condition, compared to AutoSpec’s reported 31. One main reason is that this benchmark contains recursive function calls and system/library calls which our current approach cannot handle.
For aerospace embedded software involved with complex branches, pointers and loops, we generate RSPs for 5 out of 6 selected functions of Sun Search software, demonstrating our capability; The single failed function, whose loop invariants require customized predicates, poses great challenges for LLMs. Our approach’s current limitation includes the inablity to handle recursive functions and data structures, and some constraints stemming from our selected tools (e.g. VST-A for complex arithmetic, LLM for overly complex invariants, and Frama-C for function pointers).
VII-C Summary Generation and Generalization on Motivating Example
This section demonstrates our tool’s ability to generate and generalize function summaries for the motivating example presented in Fig. 1. TripleAbsMaxFun is a C function that operates on a TripleAbsMax struct pointer, which calculates a threshold, determines the maximum absolute value of three elements within the struct, and conditionally stores this maximum.
First of all, we apply our tool to derive the RSP for TripleAbsMaxFun, the disjunction of formulas covering all 64 execution paths and capturing the function’s complete behavior. For clarity and page limit, the full summary of the function is presented in supplemental material.
Then, we apply our approach to generate from this RSP more abstract summaries within given DSLs. We propose four different DSLs for this motivating example to meet different demands, detailed as follows.
DSL for non-negative constraints
[…]
DSL for Equality Constraints
[…]
$\begin{array}[]{llll}\texttt{pIp\_fabs\_0}==\texttt{old\_pIp\_fabs\_0}\wedge\\ \texttt{pIp\_fabs\_1}==\texttt{old\_pIp\_fabs\_1}\wedge\\ \texttt{pIp\_fabs\_2}==\texttt{old\_pIp\_fabs\_2}\wedge\\ \texttt{pIp\_ret}==\texttt{old\_pIp\_ret}\\ \end{array}$
indicating that the fabs array elements do not change throughout execution.
DSL for Greater-Than-or-Equal-To Constraints
[…]
$\texttt{pIp\_fabs\_0}\geq\texttt{old\_pIp\_fabs\_0}\wedge\\  \texttt{pIp\_fabs\_1}\geq\texttt{old\_pIp\_fabs\_1}\wedge\\  \texttt{pIp\_fabs\_2}\geq\texttt{old\_pIp\_fabs\_2}\wedge\\  \texttt{pIp\_ret}\geq\texttt{old\_pIp\_ret}\wedge\\  \texttt{pIp\_tmax}\geq\texttt{old\_pIp\_fabs\_2}\wedge\\  \texttt{pIp\_tmax}\geq\texttt{old\_pIp\_fabs\_0}\wedge\\  \texttt{pIp\_tmax}\geq\texttt{old\_pIp\_fabs\_1}\end{array}$
indicating that pIp->tmax is the maximum of the absolute values of the three fabs elements, and expresses the conditional update of *pIp->ret.
DSL for Absolute Value Constraints
[…]
$\texttt{pIp\_tmax}\geq\texttt{Abs(old\_pIp\_fabs\_2)}\wedge\\    \texttt{pIp\_tmax}\geq\texttt{Abs(old\_pIp\_fabs\_0)}\wedge\\    \texttt{pIp\_tmax}\geq\texttt{Abs(old\_pIp\_fabs\_1)}\end{array}$
indicating that pIp->tmax is always greater than or equal to the absolute value of each original fabs element, reinforcing its role as the maximum absolute value.
VIII Related Work
Automated specification/summary synthesis. The automated synthesis of specifications/summaries is particularly challenging, especially with existence of loops. Traditional techniques can be classified into two categories: dynamic analysis and static analysis. Dynamic analysis infers specifications and invariants by leveraging runtime data obtained from test cases. Daikon [12] collects runtime data by executing test units and generates a candidate set of invariants based on a library of invariant templates. Sharma et al[13] adopts a similar dynamic analysis approach but furthermore checks soundness of the candidate invariants through SMT solvers. DIG [14] integrates concepts from linear algebra, geometry, and formal methods to infer invariants directly from observed program traces without need to filter out predefined templates. However, the approach based on dynamic analysis heavily depends on the quality of the test units, and an excessive number of templates can also lead to a decrease in performance.
Static analysis eliminates the dependency on test cases and explores a variety of techniques, including abstract interpretation [15] [16], liquid type inference [17] [18], abductive inference [8], predicate abstraction [19], interpolant [20, 21], symbolic execution [22] [23], and so on. Cousot et al[16] utilizes convex polyhedra along with transformations and widening operations to discover linear invariants based on abstract interpretation. Hashimoto et al[18] generate refinement types encoding maximally-preferred invariants by reducing type inference to an optimization problem. Jung et al[19] use the CDNF exact learning algorithm [24] with teacher-provided queries for invariant inference. Lin et al[20] synthesize inductive invariants using Craig interpolants derived from reachability analyses, and nonlinear Craig interpolant generation was first studied in [21]. Chalupa and Strejcek [22] propose an extension of backward symbolic execution (BSE) called loop folding to derive loop invariants that are sufficient to prove the unreachability of error locations [25]. SymInfer [23] infers program invariants from the symbolic states generated by symbolic execution of Java programs. Wu et al[26] solve the strong and weak invariant synthesis problems of polynomial programs by combining invariant templates, semidefinite programming and constraint solving techniques.
Specification/summary synthesis with machine learning. Recent advancements in machine learning [27, 28, 29, 10, 30, 31], particularly deep learning and reinforcement learning, have spurred new approaches for invariant and specification generation. CODE2INV [10], based on Counterexample-Guided Inductive Synthesis (CEGIS) [32], utilizes graph neural networks to learn neural representations of program structures. While CODE2INV employs an end-to-end reinforcement learning approach, it often faces challenges with limited performance and sparse rewards. To address these issues, LIPuS [30] establishes a clearer division of labor between the capabilities of RL and SMT solvers and introduces a novel two-dimensional reward mechanism to alleviate the sparse reward problem of CODE2INV. However, training model requires significant computational power and large datasets.
Specification/summary synthesis with LLM. Pei et al. [33] demonstrates that fine-tuning pre-trained LLMs on code can effectively predict program invariants, highlighting their potential in this field. ACInv [34] utilizes static analysis to extract essential information about loops and data structures, which is then embedded into prompts for an LLM to generate candidate invariants and abstract predicates for data structures. Loopy [35] uses Houdini algorithm to efficiently find a valid subset of inductive invariants from multiple LLM-generated candidate sets. AutoSpec [1] decomposes programs into smaller parts and uses carefully designed prompts including few-shot examples, to guide an LLM in generating specifications and verify them using Frama-C.
Abstract summary generation. Synthesizing easily understandable abstract specifications or summaries for functions has also been studied. AMURTH [36] starts from user-provided concrete semantics of operations, and employs a dual Counterexample-Guided Inductive Synthesis (CEGIS) approach to generate abstract transformers expressed within a DSL in abstract domain. SPYRO [2] expands the problem domain from specific abstract transformers to more general specification synthesis, and enhances efficiency and quality through algorithmic optimization.
Unlike existing approaches, our method uniquely integrates symbolic execution, LLMs, and formal verification in a unified framework to efficiently synthesize function summaries at multiple abstraction levels.
IX Conclusion
In this paper, we have presented a novel, end-to-end framework for automatically synthesizing multi-granularity function summaries for C programs by tightly integrating path-sensitive symbolic execution, large language models, and formal verification. Our approach generates formally-verified relatively strongest postconditions (RSPs), infers precise loop invariants via template-guided LLM synthesis and iterative refinement, and automatically abstracts these RSPs into domain-specific summaries. Experimental results on standard benchmarks and real-world aerospace code show that our tool outperforms existing methods in both precision and efficiency, achieving strong results on invariant generation, particularly on SyGuS numeric single-program benchmarks, and summarizing complex legacy functions.
For future work, we will extend our approach to handle recursive functions and recursive data structures, and devlelop richer domain-specific languages that incorporate user-defined abstract predicates. We also plan to improve our invariant generation methods to deal with these complex cases through LLM fine-tuning and specialized training procedures.